# Weyl symmetry without the traceless condition

### v3

@@ Abstract
We show that the requirement that the trace of the stress-energy tensor of matter must vanish if invariance under Weyl transformations is a symmetry of a given gravitational theory is not universal. This requirement holds whenever the masses of timelike fields are constant parameters that are not transformed by conformal transformations, or when the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-3}\rho$. In contrast, if the masses of timelike fields are point-dependent quantities transforming under conformal transformations as $m\rightarrow\Omega^{-1}m$, and the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-4}\rho$, the Weyl symmetry does not require the vanishing of the trace of the matter SET. In consequence, any matter fields, regardless of whether the trace of their stress-energy tensor vanishes or not, can be coupled to gravity. The phenomenological and physical consequences of the novel result are drawn.
@@ I Introduction
where $\Omega=\Omega(x)$ is the conformal factor ($\Omega(x)>0$), $\Phi_{i}$ are other gravitational and matter fields present, and $w_{i}$ are the conformal weights of these fields. Here we call this “Dicke’s principle.”
The high-energy physics (HEP) argument that, due to the Planck mass $M_{\text{pl}}$, a dimensionful parameter, gravity cannot be conformally invariant, seems to oppose Dicke’s principle dicke-1962 . In general, the presence of any mass term is believed to break the conformal invariance deser-1970 ; callan_prd_1970 . However, strictly speaking, this is true only if we consider constant masses that do not transform under (1): $m\rightarrow m$. As stated in bekenstein_1980 , we must differentiate between LSTs:
@@
where, in addition to CTs (1), the following transformation property of the masses is assumed:
$m\rightarrow\hat{m}=\Omega^{-1}m.$ (4)
That is, the mass parameters are fields with conformal weight $w_{m}=-1$ hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 .
According to bekenstein_1980 , scale transformations (2) are not units’ transformations; they are active enlargements of a system whose usefulness depends on the absence of a length scale. Meanwhile, WTs are units’ transformations (3) whose physical and geometrical meanings follow from the existence of some scale of length.
In the present document, as is customary when working within the framework of scalar-tensor theories (STTs) dicke-1962 ; bekenstein_1980 ; faraoni_rev ; fujii_book ; faraoni_book ; faraoni_prd_2007 ; quiros_ijmpd_2019 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 , we consider WTs (3), instead of LSTs (2). In consequence, it is assumed that the masses are point-dependent quantities that, under (1), transform like (4). Whether the mass parameter transforms or not under the Weyl transformations plays a central role in the demonstration of conformal form-invariance of the action of fundamental matter fields quiros_prd_2025 :
$S_{m}=\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (5)
@@
For fermions and gauge bosons, it has been shown in quiros_prd_2025 that if their masses are point-dependent quantities transforming as (4), the matter action (5) is form-invariant under WTs
$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\chi_{i}\rightarrow\hat{\chi}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\Omega^{-1}m,$ (7)
where $w_{\chi_{i}}$ is the conformal weight of the matter field $\chi_{i}$. But if the masses are constant parameters, then (5) is not form-invariant under (2). Take, for illustration, the Lagrangian density of the Dirac fermion: [^1] Here we use the following signature of the metric: (,+,+,+)(-,+,+,+). We follow the notation of quiros_prd_2025 (see also the Appendix A).
@@
A large body of applications of gravitational theories is associated with background matter in the form of a perfect fluid, rather than with fundamental fields. [^3] For a perfect fluid with energy density $\rho$ and barotropic pressure $p$, the SET is given by $T^{(m)}_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu}$, where $u^{\mu}=dx^{\mu}/d\tau$ is the four-velocity of the fluid ($d\tau=ids$ $\Rightarrow d\tau^{2}=-ds^{2}$). Therefore, the trace of the SET of the perfect fluid reads: $T^{(m)}=-\rho+3p$. It vanishes only for radiation with the equation of state (EOS) $p=\rho/3$. Therefore, it is of interest to look for the transformation properties of the action of perfect fluids under Weyl transformations (3). In quiros_prd_2025 , it has also been shown that, in addition to the action of fundamental matter fields, that of perfect fluids is form-invariant under WTs as well. Actually, the action of a perfect fluid with energy density $\rho$ and barotropic pressure $p$ can be written in either form: $S_{\text{fluid}}=-\int d^{4}x\sqrt{-g}\,\rho$ or $S_{\text{fluid}}=\int d^{4}x\sqrt{-g}\,p$. Unless the perfect fluid couples explicitly to the curvature (which is not the case in this paper), both Lagrangian densities are equivalent faraoni_2009 . Here, for definiteness, we choose the former action for the perfect fluid. According to dimensional arguments, if we assume point-dependent masses that transform as $m\rightarrow\Omega^{-1}m$ under CTs (1), the energy density of the fluid transforms as $\rho\rightarrow\Omega^{-4}\rho$. In fact, the energy density has units: $M]/[L]^{3}$, where $M]\rightarrow\Omega^{-1}[M]$ and $L]\rightarrow\Omega[L]$ are the transformation laws of the units of mass and length under CTs, respectively. Hence, $M]/[L]^{3}\rightarrow\Omega^{-4}[M]/[L]^{3}$. The barotropic pressure of the fluid $p$ transforms in the same way: $p\rightarrow\Omega^{-4}p$. Since under (1), $\sqrt{-g}\rightarrow\sqrt{-\hat{g}}=\Omega^{4}\sqrt{-g}$ and $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, then $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\hat{\rho}$. That is, the action of the perfect fluid is form-invariant under WTs. Note, however, that if we assume that the mass unit is not transformed, then, under CTs $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$, which means that the action of the perfect fluid is not form-invariant: $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\Omega\hat{\rho}$.
@@
The requirement of vanishing of the matter SET trace, in connection with LSS in curved spacetime, was derived in deser-1970 , in the case of a scalar field non-minimally coupled to gravity. [^4] The vanishing of the matter SET trace as a requirement for conformal symmetry to hold ccj_ann_phys_1970 was primarily investigated within the HEP framework, where gravity is ignored, so the background spacetime is flat space with the Minkowski metric. Note, in this regard, that, within the HEP context, conformal symmetry refers to the 15-parameter conformal group that consists, in addition to the Poincaré group, of the 1-parameter group of dilatations and the 4-parameter special conformal group kastrup_1966 ; jackiw_1972 . These are spacetime transformations. In contrast, the conformal transformations that we consider in this paper are transformations of fields, including the spacetime metric and any matter fields that could be coupled to gravity, that do not affect either the spacetime points or the coordinates that label these points. That is, transformations (1) are not spacetime diffeomorphisms. For that reason, in this document, to avoid confusion between conformal transformations (1) and the HEP 15-parameter group of transformations, we will refer to the former mostly as Weyl transformations. In consequence, we say, for example, “Weyl symmetry” instead of “conformal symmetry.” The gravitational action of the theory investigated in this bibliographic reference, called here “conformal general relativity” (CGR), reads:
@@
where $R$ is the curvature scalar, $\phi$ is the non-minimally coupled scalar field with mass dimensions, and we use the following shorthand notation: $(\partial\phi)^{2}\equiv g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$. In deser-1970 , it is demonstrated that the form-invariance of (9) under LSTs (2), is possible thanks to the vanishing trace of the matter SET: $T^{(m)}\equiv g^{\mu\nu}T^{(m)}_{\mu\nu}=0$.
@@
In this document, we shall demonstrate that the traceless condition $T^{(m)}=0$ is not a universal requirement for Weyl symmetry to materialize. The result of reference deser-1970 , that only matter fields with traceless SET can be coupled to gravity in CGR theory (9), is right only if we assume that the masses of timelike fields are constant parameters that are not transformed by CTs (1), so the energy density of a perfect fluid transforms as: $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$; that is, if we consider local scale transformations (2) instead of Weyl transformations (3). Under this assumption, the Lagrangian density of the matter fields breaks LSS. In contrast, if we assume that the masses of timelike fields are point-dependent quantities that, under the CTs (1), transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, and consequently the energy density of the perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, the Lagrangian density of the fundamental matter fields, as well as the Lagrangian density of perfect fluids, are form-invariant under the WTs (3):
$\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,\Omega^{-4}\,L_{m}(\Omega^{-w_{\chi_{i}}}\hat{\chi}_{i},\Omega^{-2}\hat{g}_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{g}_{\mu\nu}),$
that is,
$\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{g}_{\mu\nu}).$ (10)
Similarly, the action of a perfect fluid is form-invariant under (7) as well: $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\hat{\rho}$. In these cases, since the Lagrangian density of matter is form-invariant under WTs, the Ward identity associated with Weyl invariance leads to the following nonvanishing variational derivative: $\delta{\cal L}_{m}/\delta\phi=\sqrt{-g}\,T^{(m)}/\phi$, which provides the term with the trace of the stress-energy tensor of matter in the EOM of the scalar field $\phi$. In contrast, if the Lagrangian density of the matter fields is not form-invariant under CTs (1), the Ward identity does not hold and $\delta{\cal L}_{m}/\delta\phi=0$.
@@
Our main hypothesis in this work is that Dicke’s principle holds in nature. That is, the laws of gravity are form-invariant under Weyl transformations (3). In consequence, we consider that the masses of fundamental timelike matter fields $m_{i}$ are point-dependent quantities bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 : $m_{i}(x)=\kappa_{i}\,\phi(x)$, where $\kappa_{i}$ are dimensionless constants. [^5] This choice is dictated by the fact that, since the scalar field $\phi$ has mass dimensions, under WTs the mass parameter transforms as $\phi$ does; that is, $m\propto\phi$. These transform as (4) under Weyl transformations, while the energy density of a perfect fluid transforms as $\rho\rightarrow\Omega^{-4}\rho$.
@@
We have organized the present document as follows. In Section II, we set up the notation and briefly review the most important aspects of the work of reference deser-1970 , where the invariance of CGR under LSTs (2) is considered. This is required for further comparison with the results of Section III, where the equations of motion (EOM) are derived from the overall CGR action under the assumptions of the present work, above all form-invariance of the matter action under WTs (3). In Section IV, we derive some physical consequences of the CGR theory, in particular, the fifth force required by form-invariance under WTs, which does not interact with radiation, thus acting as a ”dark force.” In this section, we also discuss the two possible approaches to CTs in the configuration space: i) passive and ii) active approaches. We shall show that only the active approach is suitable for the investigation of the physical and phenomenological consequences of the Weyl symmetry. In Section V, we discuss gauge freedom within CGR, in connection with Weyl symmetry, under the active approach to WTs. We draw the similitude of the picture arising from gauge freedom within the CGR theory and the “many-worlds” interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 . In Sections VI and VII, we expose some important cosmological consequences of the present setup, including the possible explanation of the galactic rotation curves in some gauge of the CGR theory, as well as some encouraging aspects of Weyl invariant cosmology. Possible observational consequences of CGR theory, mainly related to measurements of the redshift of frequencies, are analyzed in Section VIII. The novel results of the present work are discussed in Section IX, accompanied by brief concluding remarks.

### v4

# Could an overlooked symmetry of the classical laws of gravity explain the cosmological puzzles?

@@ Abstract
We show that if the masses of timelike fields are point-dependent quantities transforming under conformal transformations as $m\rightarrow\Omega^{-1}m$, and the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-4}\rho$, form-invariance under Weyl transformations is an actual symmetry of the gravitational interactions of matter. That is, under the mentioned circumstances, Weyl symmetry allows any matter field to be coupled to gravity. The phenomenological and physical consequences of the novel result, including the “many worlds” interpretation of gauge freedom, are drawn. We explore, in particular, a possible explanation of two major cosmological puzzles: dark matter and dark energy, as a consequence of Weyl symmetry. Quantum-mechanical removal of the spacetime singularities in this framework is briefly discussed as well.
@@ I Introduction
where $\Omega=\Omega(x)$ is the conformal factor ($\Omega(x)>0$), $\Phi_{i}$ are other gravitational and matter fields present, and $w_{i}$ are the conformal weights of these fields. Here we call this invariance “Dicke’s principle.”
The high-energy physics (HEP) argument that, due to the Planck mass $M_{\text{pl}}$, being a dimensionful parameter, gravity cannot be conformally invariant, seems to oppose Dicke’s principle dicke-1962 . In general, the presence of any mass term is believed to break the conformal invariance deser-1970 ; callan_prd_1970 . However, strictly speaking, this is true only if we consider that constant masses do not transform under (1): $m\rightarrow m$. As stated in bekenstein_1980 , we must differentiate between LSTs:
@@
where, in addition to CTs (1), the transformation of the masses $m\rightarrow\Omega^{-1}m$, is assumed. That is, the mass parameters are fields with conformal weight $w_{m}=-1$ hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 .
According to bekenstein_1980 , scale transformations (2) are not unit transformations; they are active enlargements of a system whose usefulness depends on the absence of a length scale. Meanwhile, WTs (3) are units’ transformations whose physical and geometrical meanings follow from the existence of some scale of length.
In the present document, as is customary within the framework of scalar-tensor theories (STTs) dicke-1962 ; bekenstein_1980 ; faraoni_rev ; fujii_book ; faraoni_book ; faraoni_prd_2007 ; quiros_ijmpd_2019 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 , we consider WTs (3), instead of LSTs (2). In consequence, it is assumed that the masses are point-dependent fields that, under (1), transform like in (3).
Whether the mass parameter transforms or not under CTs plays a central role in the demonstration of conformal form-invariance of the action of fundamental matter fields quiros_prd_2025 :
$S_{m}=\int d^{4}x\,{\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (4)
@@
For fermions and gauge bosons, it has been shown in quiros_prd_2025 that if their masses are point-dependent quantities transforming as in (3), the matter action (4) is form-invariant under WTs
$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\chi_{i}\rightarrow\hat{\chi}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\Omega^{-1}m.$ (6)
But if the masses are constant parameters, then (4) is not form-invariant under (2). Take, for illustration, the Lagrangian density of the Dirac fermion: [^1] Here we use the following signature of the metric: (,+,+,+)(-,+,+,+). We follow the notation of quiros_prd_2025 .
@@
A large body of applications of gravitational theories is associated with background matter in the form of a perfect fluid, rather than with fundamental fields. Therefore, it is of interest to investigate the transformation properties of the action of perfect fluids under Weyl transformations (3). In quiros_prd_2025 , it has also been shown that, in addition to the action of fundamental matter fields, that of perfect fluids is form-invariant under WTs as well. Actually, the action of a perfect fluid with energy density $\rho$ and barotropic pressure $p$ can be written in either form: $S_{\text{fluid}}=-\int d^{4}x\sqrt{-g}\,\rho$ or $S_{\text{fluid}}=\int d^{4}x\sqrt{-g}\,p$. Unless the perfect fluid couples explicitly to the curvature (which is not the case in this paper), both Lagrangian densities are equivalent faraoni_2009 . Here, for definiteness, we choose the former action for the perfect fluid. According to dimensional arguments, if we assume point-dependent masses that transform as $m\rightarrow\Omega^{-1}m$ under CTs (1), the energy density of the fluid transforms as $\rho\rightarrow\Omega^{-4}\rho$. In fact, the energy density has units: $M]/[L]^{3}$, where $M]\rightarrow\Omega^{-1}[M]$ and $L]\rightarrow\Omega[L]$ are the transformation laws of the units of mass and length under CTs, respectively. Hence, $M]/[L]^{3}\rightarrow\Omega^{-4}[M]/[L]^{3}$. The barotropic pressure of the fluid $p$ transforms in the same way: $p\rightarrow\Omega^{-4}p$. Since under (1), $\sqrt{-g}\rightarrow\sqrt{-\hat{g}}=\Omega^{4}\sqrt{-g}$ and $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, then $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\hat{\rho}$. That is, the action of the perfect fluid is form-invariant under WTs. Note, however, that if we assume that the mass unit is not transformed under CTs, then $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$, which means that the action of the perfect fluid is not form-invariant: $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\,\Omega\sqrt{-\hat{g}}\hat{\rho}$.
@@
The requirement of vanishing of the matter SET trace, in connection with local scale symmetry (LSS) in curved spacetime, was derived in deser-1970 , in the case of a scalar field non-minimally coupled to gravity. [^3] The vanishing of the matter SET trace as a requirement for conformal symmetry to hold ccj_ann_phys_1970 was primarily investigated within the HEP framework, where gravity is ignored, so the background spacetime is flat space with the Minkowski metric. Note, in this regard, that, within the HEP context, conformal symmetry refers to the 15-parameter conformal group that consists, in addition to the Poincaré group, of the 1-parameter group of dilatations and the 4-parameter special conformal group kastrup_1966 ; jackiw_1972 . These are spacetime transformations. In contrast, the conformal transformations that we consider in this paper are transformations of fields, including the spacetime metric and any matter fields that could be coupled to gravity, that do not affect either the spacetime points or the coordinates that label these points. That is, transformations (1) are not spacetime diffeomorphisms. For that reason, in this document, to avoid confusion between conformal transformations (1) and the HEP 15-parameter group of transformations, we will refer to the former mostly as Weyl transformations. In consequence, we say, for example, “Weyl symmetry” instead of “conformal symmetry.” The gravitational action of the theory investigated in this bibliographic reference, called here “conformal general relativity” (CGR), reads:
@@
where $R$ is the curvature scalar, $\phi$ is the non-minimally coupled scalar field with mass dimensions, and we use the following shorthand notation: $(\partial\phi)^{2}\equiv g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$. In deser-1970 , it is demonstrated that the form-invariance of (8) plus a matter action piece $S_{m}=\int d^{4}x\sqrt{-g}\,L_{m}$ under LSTs (2), is possible only if the trace of the matter SET vanishes: $T^{(m)}\equiv g^{\mu\nu}T^{(m)}_{\mu\nu}=0$. Here, we shall demonstrate that the traceless condition $T^{(m)}=0$ is not a universal requirement for Weyl symmetry to materialize. The result of reference deser-1970 , that only matter fields with traceless SET can be coupled to gravity in CGR theory (8), is correct only if we assume that the masses of timelike fields are constant parameters that are not transformed by CTs (1), so the energy density of a perfect fluid transforms as: $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$; that is, if we consider local scale transformations (2) instead of Weyl transformations (3). Under this assumption, the Lagrangian density of the matter fields breaks LSS. In contrast, if we assume that the masses of timelike fields are point-dependent fields that, under the CTs (1), transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, and consequently the energy density of the perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, the Lagrangian density of the fundamental matter fields, as well as the Lagrangian density of perfect fluids, are form-invariant under the WTs (3):
$\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{g}_{\mu\nu}),$ (9)
and $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\hat{\rho}$, respectively.
@@
Our primary hypothesis in this work is that Dicke’s principle holds in nature. That is, the laws of gravity are form-invariant under Weyl transformations (3). In consequence, we consider that the masses of the fundamental timelike matter fields $m_{i}$ are point-dependent fields bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 : $m_{i}(x)=\kappa_{i}\,\phi(x)$, where $\kappa_{i}$ are dimensionless constants. [^4] This choice is dictated by the fact that, since the scalar field $\phi$ has mass dimensions, under CTs the mass parameter transforms as $\phi$ does; that is, $m\propto\phi$. These transform as in (3) under Weyl transformations, while the energy density of a perfect fluid transforms as $\rho\rightarrow\Omega^{-4}\rho$. In these cases, since the Lagrangian density of matter is form-invariant under WTs, the Ward identity associated with Weyl invariance leads to the following nonvanishing variational derivative: [^5] In contrast, if the mass parameter $m$ is an untransformed constant, the Lagrangian density of the matter fields is not form-invariant under CTs (1), the Ward identity does not hold and $\delta{\cal L}_{m}/\delta\phi=0$. $\delta{\cal L}_{m}/\delta\phi=\sqrt{-g}\,T^{(m)}/\phi$, which provides the term with the trace of the stress-energy tensor of matter in the EOM of the scalar field $\phi$.
We will aim to investigate the phenomenological and physical consequences of the resulting Weyl-invariant theory of gravity, that is, of CGR. In particular, the possibility of finding an alternative explanation for two major cosmological puzzles: dark matter (DM) and dark energy (DE), will be explored.
@@
We have organized the present document as follows. In Section II, we set up the notation and briefly review the most important aspects of the work of reference deser-1970 , where the invariance of CGR under LSTs (2) is considered. This is required for further comparison with the results of Section III, where the equations of motion (EOM) are derived from the overall CGR action under the assumptions of the present work, above all form-invariance of the matter action under WTs (3). In this Section, we also derive some physical consequences of the CGR theory, in particular, the fifth force required by form-invariance under WTs, which does not interact with radiation, thus acting as a ”dark force” with the potential to explain the “dark sector” of present cosmology. In Section IV, we discuss the two possible approaches to CTs in the configuration space quiros_prd_2025 : i) passive and ii) active approaches. We shall show that only the active approach is suitable for the investigation of the physical and phenomenological consequences of the Weyl symmetry. In Section V, we discuss gauge freedom within CGR, in connection with Weyl symmetry, under the active approach to WTs. We draw the similitude of the picture arising from gauge freedom within the CGR theory and the “many-worlds” interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 . In Sections VI and VII, we expose some important cosmological consequences of the present setup, including the possible explanation of the galactic rotation curves in some gauge of the CGR theory, as well as some encouraging aspects of Weyl-invariant cosmology. Possible observational consequences of CGR theory related to frequency redshift measurements are analyzed in Section VIII, where a natural explanation of the measured redshifts of supernovae of type Ia (SNIa) is provided without requiring dark energy. A discussion of quantum-mechanical removal of spacetime singularities in the present many-worlds framework is given in Section IX. This is followed by brief concluding remarks of the present work in Section X. An appendix A is included where the curvature redshift is derived as a solution of the null-geodesic equation.